s=3t+1
find the range of the function if the domain is {-3,0,3}

Answer by AnlytcPhil(1810)

s = 3t + 1 
find the range of the function if 
the domain is {-3,0,3}

Plug in the -3 for t.

s = 3t + 1
s = 3(-3) + 1
s = -9 + 1
s = -8

So -8 is one member of the range. 

Plug in the 0 for t.

s = 3t + 1
s = 3(0) + 1
s = 0 + 1
s = 1

So 1 is another member of the range.

Plug in the 3 for t.

s = 3t + 1
s = 3(3) + 1
s = 9 + 1
s = 10

So 10 is the last member of the range. 

So the range is {-8, 1, 10}

That wasn't very hard, was it?

Edwin